# Better Together? A Field Experiment on

### v2

@@ 2.3 Validity of the Algorithm Tool
We assess the usefulness of the algorithmic tool using proprietary state hospital records that were not available to the algorithm developers (Figure A.3). Subfigure (i) plots the relationship between algorithm risk quintile and a standardized aggregate index of potential maltreatment-related hospitalizations that is used as a primary outcome (described in Section 3.5). A greater algorithm risk score predicts higher levels of child harm in the study’s control group sample. Likewise, other subfigures of Figure A.3 present a robust, monotonic relationship between the tool’s risk score and other indicators of child maltreatment risk such as child removals, re-referrals, and investigator ratings of risk conditional on visiting a family. The patterns are consistent across the primary race-ethnic, gender, and socioeconomic status groups used for heterogeneity analyses. These findings confirm that the algorithm provided workers with a signal of future child outcomes.
Figure A.4 provides an illustrative example of the potential utility of an algorithm tool. The figure again shows that the tool is predictive of ground-truth medical risk, an outcome that the algorithm was not directly trained to predict. An above-median algorithm risk score is associated with a greater incidence of emergency room (ER) visits – a signal of potential abuse or neglect – consistently over time. Around the time that children are called into CPS, however, ER visits are elevated among lower-risk children to a level comparable to that of the high-risk group. Therefore, although the tool predicts potential harm, it is ex-ante possible that in some instances CPS agencies could struggle to discern underlying risk at the moment of contact, when both lower- and higher-risk children look similar on some dimensions. By providing a summary statistic about underlying household circumstances, it is plausible that access to an algorithmic tool could improve decision quality and child outcomes.
@@ 3.1 Data Sources
Statewide hospital inpatient and emergency room (ER) records were linked through a partnership with the state’s hospital association, spanning January 2020 through June 2022.999 Records were matched probabilistically using person-level identifiers (full name, date of birth, and gender) and include fields for: dates of admission and discharge, point of origin, facility type, admission priority, charged amount, diagnosis codes, and procedure codes. Match probabilities were 95 percent or greater for nearly all children. Notably, we requested information on all hospitalizations in the state – not just hospitalizations in the trial county – to address concerns about differential migration. 86 percent of matched hospitalizations occurred within the trial county, and 93 percent occurred within either the trial county or a neighboring county.
@@ 3.2 Sample Characteristics
Table 1 presents descriptive statistics for the full analysis sample in column 1, and a hospital outcomes sample in column 2. Column 2 consists of referrals within the first year of the trial, for whom we observe several months of potential hospitalization following randomization. The reasons for focusing on this subsample for health outcomes is to allow sufficient time for outcomes to appear and to address post-trial contamination of treatment status, which would both bias our estimates towards zero. [^10] Appendix C transparently enumerates all deviations from our pre-analysis plan, and clarifies the reasons for doing so. We had not pre-specified a subsample for hospital outcomes, but our results are robust to using the full sample of observations (Table A.4). Indeed, at the completion of the trial on March 30th 2022, the algorithm was made accessible for all incoming calls, including for children who were previously assigned to the control group. Workers could then observe algorithm scores for all children, attenuating potential treatment effects. Restricting to a subset of children referred during the first year of the trial allows us to better estimate the effects of a world with the algorithmic tool compared to a world without it.
The full sample presented in column 1 of Table 1 is evenly split in terms of gender, only two thirds of children are recorded as white (including Hispanic ethnicity) in the data system, 3.6 percent as Black (including Hispanic ethnicity), and race is unknown for most of the remaining children. As is true of many child protection agencies, Black and Hispanic children are overrepresented in the sample relative to their share of county population. Although ethnicity is unknown for 34 percent of children, at least 17 percent of children in the sample are Hispanic.
Table 1: Sample Characteristics
(1) (2)
Full Sample Hospital Outcomes Sample
Referred in First Year of Trial 0.73 1
Panel A: Child Demographics
Female 0.52 0.51
Age at Initial Randomization 10.5 10.4
White 0.66 0.69
Black 0.036 0.041
Missing Race 0.29 0.25
Hispanic 0.17 0.18
Missing Ethnicity 0.34 0.30
Panel B: Referral Details
N Children Listed on Referral 2.24 2.26
Mandated Reporter 0.74 0.73
Neglect Allegation 0.77 0.77
Abuse Allegation 0.43 0.43
Panel C: Child Experiences
Total Referrals in Trial 2.19 2.42
Any Re-referral in Trial 0.42 0.48
Removed from Home during Trial 0.022 0.027
Any Hospital Visit After Referral 0.42 0.47
N Hospital Visits After Referral 1.64 1.95
N Unique Children 4,681 3,431
N Referrals 2,832 2,068
N Family Clusters 2,810 2,055
Notes: This table describes the characteristics of children who were reported to CPS and successfully randomized during the trial. Column 1 describes the full sample. Column 2 describes the subset of children referred to CPS during the first year of the trial, for whom we observe at least several months of possible hospital visits after randomization.
A child’s referral has on average 2.24 (median of 2) children listed. Some referred families include only a single newborn child, whereas others consist of several older children. 74 percent of calls originate from mandated reporters, including school employees, physicians, and law enforcement. Referrals are tagged at the hotline with allegations from up to 24 categories. “Injurious environment,” a form of neglect, is the most common allegation recorded by the agency, appearing for 62.1 percent of children. The two other most prominent allegation categories are physical abuse (20.3 percent) and parental substance abuse (14.7 percent). Overall, three quarters of referrals have at least one allegation of neglect, and 43 percent have at least one allegation of abuse (Table 1, Panel B).
The hospital outcomes subsample (Table 1, Column 2), which includes all referrals during the first year of the trial, is comparable to the full sample in terms of observable characteristics. By construction, however, the hospital outcomes subsample has greater levels of CPS contact and hospital interaction following a child’s initial referral in the trial. 48 percent of children in the hospital outcomes sample are re-referred to CPS (42 percent in the full sample), 47 percent are hospitalized after their initial referral (42 percent in the full sample), and the mean number of hospital visits is 1.95 (1.64 in the full sample).
@@ 3.4 Empirical Strategy
where $Y_{i}$ is the outcome of interest and $\epsilon_{i}$ an error term. Due to an unintentional coding error from the implementing partner, a small subset of observations were randomized only conditional on sibling group size $s(i)$, leading us to include corrective “randomization controls” $\gamma_{s(i)}$ in our specifications (see Appendix C for more details). The key variable, $AlgoAccess_{r(i)}$, is a binary indicator equal to one if the algorithm tool was available to the team during their discussion of the child’s first referral $r(i)$ during the trial, and zero if the algorithm was not available. Standard errors are clustered conservatively at the household (overlapping family network) level. [^12] A household cluster is defined conservatively as any extended network of individuals who ever appeared jointly on a referral. In practice, there is no difference from clustering at the referral level (Appendix Table A.4).
@@
The effectiveness of algorithmic decision aids depends on if and when workers choose to use them. Since workers were strongly encouraged but not required to consult the tool in our setting, the policy-relevant treatment effect is the intent-to-treat (ITT) effect of access to the algorithm. Using de-identified discussion notes, we ascertain that workers wrote down algorithm scores in at least 73 percent of instances when available in our sample (Table A.2). This fraction of observed and recorded scores is relatively constant across the algorithm risk score distribution (Table A.2, column 2). The recorded share is likely an underestimate: workers may have consulted the tool but forgotten to write down the score in their meeting notes. Our estimates show that workers regularly chose to consult the algorithm, even though there were no professional repercussions for not using it.
Random assignment is important for estimating unbiased causal effects of algorithm availability on outcomes. We check for evidence of random assignment with a balance test, regressing assignment status on a vector of salient pre-randomization observable characteristics. Table A.3 suggests that the algorithm was randomized successfully: we do not find joint significance of the coefficients on observable characteristics. The p-value of the F-test is 0.503 in the full sample, and 0.106 in the hospital sample. Our main findings are robust to controlling for all the variables in the balance test (Row "With Controls" of Appendix Table A.4), confirming that our results cannot be explained by differences in an extensive set of observables.
@@ 3.5 Outcomes: Proxies for Child Harm
We measure effects on child wellbeing using a pre-specified set of hospitalization outcomes across the entire state. These include: a child’s number of high-priority hospital admissions (emergency, urgent, or trauma-related), admissions with injury ICD (diagnosis) codes, preventable emergency room visits, visits with ICD codes for intentional harm (assault or self-harm), and visits with ICD codes for maltreatment, which are rare. Preventable ER visits are defined as the number of child admissions for ambulatory care sensitive conditions (ACSC, classified using Appendix 8 of Carey et al. 2017). For example, some ER visits for asthma may have been avoidable if a child had received proper preventative treatment and supervision.
@@
These categories of hospitalization have been clinically and empirically associated with child maltreatment. Maltreatment ICD codes are a measure of physician-confirmed maltreatment, but notoriously underestimate overall child maltreatment (Scott et al., 2009; Schnitzer et al., 2011). Consequently, researchers have primarily used and validated injuries as a leading proxy for maltreatment (Schnitzer et al., 2011; Vaithianathan et al., 2020). Prior research has also demonstrated a link between maltreatment and intentional injury, namely self-harm, which can often result from adverse childhood experiences such as parental neglect and sexual abuse(Liu et al. 2018, Vaithianathan et al. 2020). High-priority admissions are a related signal of potential acute harm, and avoidable ER visits are particularly associated with possible neglect and insufficient parental supervision. Our outcome categories therefore cover a range of maltreatment types: physical harm, psychological harm, and neglect.
@@
Our preferred outcome is an aggregated “harm index” of pre-specified hospitalization outcomes. A high Cronbach’s alpha ($\alpha=0.81$), reflecting internal consistency across the outcomes, provides some precedent and justification for combining them into a composite index. The index is the standardized average of the five variables described above, each themselves standardized. Variables are standardized (mean 0, variance 1) using the control sample, and a lower value of the harm index is considered better for the child. This index was chosen as the most standard and agnostic way to combine these outcomes, in the absence of a pre-specified index. In Appendix Table A.4 we also consider alternative indices: a harm index that uses only a binary, extensive margin for all five outcome categories (e.g., any high-priority, any injury, etc.), a first principal component, and a weighted index where weights are determined by the inverse of the sample variance-covariance matrix to accommodate dependence across outcomes (O’Brien, 1984). Per our pre-analysis plan, all outcomes are measured following 30 days after randomization to accommodate a period of potential mechanical increases in child receipt of medical care from CPS interactions, such as post-investigation medical care for previously untreated conditions. Results are robust to excluding the first 60 days or to including the first 30 days (Table A.4).
@@ 4 Results
Despite Child Protective Service’s primary objective of protecting children, few studies have produced causal estimates of the impact of CPS policies on child injury and other health outcomes. We link statewide hospitalization records from the trial state’s hospital association (Section 3.1) to measure the effect of algorithmic information on ground-truth outcomes. This section covers three main findings: a reduction in child harm, a reduction in child harm disparities, and a reduction in racial CPS surveillance disparities.
@@ 4.1 Impact of Algorithm Availability on Child Health and Wellbeing
Table 2 presents estimated effects of algorithm support on child wellbeing for the outcomes specified in the previous subsection, using equation (1). Column 1 shows that tool availability reduced a child’s future harm by 0.061 standard deviations (SDs) compared to children in the control group ($p<0.05$). This estimate is particularly striking given that there were no observable differences in the intervention and control group’s harm indices in the 180 days prior to initial referral (Appendix Table A.3; estimates are robust to controlling for hospitalization history). A 0.061 SD reduction in the harm index is sizable in this context. For comparison, children removed to foster care during the trial – a proxy for severe maltreatment or significant danger – had an accumulated harm index that was 0.34 standard deviations greater than those children who were not removed (0.34 SD difference in the control group, 0.46 SD difference for the full sample). The intent-to-treat effect from the algorithm is approximately 18 percent of this benchmark and – assuming our first stage does not underestimate tool use – the effect on compliers (-0.083 SD; Table A.4) is as large as 24 percent of the benchmark. Reassuringly, we find no effects for 12 percent of children assigned to treatment for whom workers tried but could not see the score due to temporary unavailability (Table A.5, column 1 placebo specification).
@@
(1) (2) (3) (4) (5) (6) (7)
Harm
Index
@@
Top 1% in
Harm Index
Algorithm Available -0.061** -0.10* -0.060** -0.020 -0.0014 -0.0072 -0.011***
(0.030) (0.056) (0.024) (0.022) (0.0043) (0.0047) (0.0037)
Randomization Controls Yes Yes Yes Yes Yes Yes Yes
Control Mean 0.000 0.660 0.210 0.168 0.013 0.020 0.016
Effect (%) -16 -29 -12 -11 -36 -67
Power 0.81 0.80 0.82 0.48 0.18 0.48 0.86
Observations 3,431 3,431 3,431 3,431 3,431 3,431 3,431
@@
Notes: This table presents estimates of the impact of algorithm availability on child injuries and other hospitalization outcomes. The harm index is constructed as a standardized average of the following standardized outcomes: number of high-priority admissions (emergency, urgent, and trauma), number of admissions with listed injury, number of avoidable ER visits, number of admissions with a confirmed maltreatment code, and number of visits with an intentional injury (assault, self-harm) code. A lower value of the harm index is considered better. Columns 2-6 present the effects on each of the outcomes before standardization. Column 7 estimates the effect on the likelihood of being in the top one percent of the harm index. The sample is restricted to children in the hospital outcomes sample. Randomization procedure controls are included throughout. We report effects as a percent of the control mean and the power to detect significant effects at the bottom of the table. Power is calculated using observed means and standard deviations in treatment and control groups, assuming 1,000 household clusters per group, an average cluster size of two, a coefficient of variation of 0.56, an intraclass correlation (ICC) of 0.4, and a significance level of 10 percent. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
Columns 2-6 of Table 2 present effects on each of the five outcomes that make up the harm index. Point estimates are negative across all five outcomes, and significant for all well-powered outcomes, suggesting that providing access to the algorithm may have improved outcomes on all five dimensions. Furthermore, the point estimates are large relative to the means in the control group, ranging from -11% to -36%. In particular, the algorithmic tool reduced the average number of future injury admissions by 29 percent, or 0.06 admissions relative to a mean of 0.21 ($p<0.05$).
Column 7 shows that the algorithmic tool was particularly effective at reducing extreme values of child harm ($p<0.01$): Significantly fewer children placed in the top one percent of the harm index (defined over the intervention and control groups together) when the tool was available. Accumulated harm for this top percent ranged from 2.6 to 29 standard deviations above the mean, with the median child in this group having a high priority hospital visit every three months. Relatedly, benefits of the algorithm appear to be greatest for children in the top quintile of algorithm-predicted risk scores, who may have been at greatest risk of maltreatment (Appendix Figure A.6). The majority of children with an explicit maltreatment-related hospital ICD code had a harm index in the top one percent of the harm distribution, which suggests that the algorithm may have reduced child maltreatment.
The reduction in child harm does not appear to be driven by changes in willingness to seek care. Tool access did not significantly reduce the overall incidence or frequency of hospital interactions (Columns 4 and 5 of Appendix Table A.5). The placebo outcome of child cancer-related hospitalizations (pre-specified following Vaithianathan et al. 2020) is unaffected, nor for lower-priority hospital visits that are less likely to indicate severe maltreatment. (columns 2 and 3 of Appendix Table A.5). Therefore, providing algorithm support significantly decreased the number of hospitalizations associated with maltreatment, but not the total number of hospital interactions.
Changes in other policy-relevant outcomes are broadly consistent with a reduction in child harm from access to the algorithm. First, only for children under age 10, practitioners have identified a set of ICD codes suggestive of maltreatment, with an emphasis on physical abuse(Schnitzer et al. 2011). Point estimates suggest that providing the algorithm may have reduced the incidence of practitioner-defined maltreatment visits by about 36 percent, though estimates are imprecise due to restricting the sample to children under age 10, and because of the relative rarity of hospitalizations with practitioner-identified codes (Appendix Table A.5, column 6). Second, we examine the impact of the tool on child re-referrals and removals in Appendix Table A.6. Point estimates suggest that providing the algorithm reduced children’s re-referrals to CPS (-5 percent), but estimates are again not statistically distinguishable from zero. We find no evidence of a reduction in child removals from the home, but this could be because removals are one of the means by which CPS prevents child maltreatment. find suggestive evidence of a five percent reduction in medical charges to public payer insurance for children with the algorithm, though these effects are not statistically significant (Appendix Table A.5, columns 7-8).
@@
We present a series of robustness tests for the main effects of the algorithm on child harm in Appendix Table A.4. The findings are robust to: a permutation test ($p=0.04$); to controlling for all baseline characteristics (including prior hospitalizations); excluding the first 60 days after a referral when prior injuries could be discovered as a result of a CPS-triggered health examination; including all days after a referral; and to including children referred after October 2021, for whom we have only a few months of hospital data available after the referral. The findings are also robust to other ways of combining outcomes into a harm index, such as using a weighted index, using an index with only outcomes’ extensive margins, and using a first principal component. Finally, results are robust to alternative clustering of standard errors and to a bounding exercise for treatment non-compliance. [^13] As discussed in Appendix C, $n=61$ children in the hospital sample control group had the algorithmic tool available due to changes in family structure.
Absence of Spillovers and Learning.

### v3

@@ 2.3 Validity of the Algorithm Tool
We assess the usefulness of the algorithmic tool using proprietary state hospital records that were not available to the algorithm developers (Figure A.3). Subfigure (i) plots the relationship between algorithm risk quintile and a standardized aggregate index of potential maltreatment-related hospitalizations that is used as a primary outcome (described in Section 3.5). A greater algorithm risk score predicts higher levels of child harm in the study’s control group sample. Likewise, other subfigures of Figure A.3 present a robust, monotonic relationship between the tool’s risk score and other indicators of child maltreatment risk such as child removals and investigator ratings of risk conditional on visiting a family. The patterns are consistent across the primary race-ethnic, gender, and socioeconomic status groups used for heterogeneity analyses. These findings confirm that the algorithm provided workers with a signal of future child outcomes.
@@ 3.1 Data Sources
Statewide hospital inpatient and emergency room (ER) records were linked through a partnership with the state’s hospital association, spanning January 2020 through June 2022.101010Records were matched probabilistically using person-level identifiers (full name, date of birth, and gender) and include fields for: dates of admission and discharge, point of origin, facility type, admission priority, charged amount, diagnosis codes, and procedure codes. Match probabilities were 95 percent or greater for nearly all children. Notably, we requested information on all hospitalizations in the state – not just hospitalizations in the trial county – to address concerns about differential migration. 86 percent of matched hospitalizations occurred within the trial county, and 93 percent occurred within either the trial county or a neighboring county.
@@ 3.2 Sample Characteristics
Appendix Table A.2 presents descriptive statistics for the full analysis sample. Our sample is evenly split in terms of gender. Two-thirds of children are recorded as white in the data system, 3.6 percent as Black, and race is unknown for most of the remaining children. Race and ethnicity are not mutually exclusive. Although ethnicity is unknown for 34 percent of children, at least 17 percent of children in the sample are Hispanic. As is true of many child protection agencies, Black and Hispanic children are overrepresented in the sample relative to their share of county population.
The average referral has 2.24 (median of 2) children listed. Children are referred on average 2.2 times during the trial, and two percent of children are removed. Overall, three quarters of referrals have at least one allegation of neglect, and 43 percent have at least one allegation of abuse (Appendix Table A.2, Panel B). 43 percent of children have a hospital visit after their referral, with an average of 1.7 hospital visits (Appendix Table A.2, Panel C).
@@ 3.4 Empirical Strategy
where $Y_{i}$ is the outcome of interest and $\epsilon_{i}$ an error term. Due to an unintentional coding error from the implementing partner, a small subset of observations were randomized only conditional on sibling group size $s(i)$, leading us to include corrective sibling-group-size fixed effects $\gamma_{s(i)}$, hereafter called “randomization controls,” in our specifications (see Appendix C for more details). The key variable, $AlgoAccess_{r(i)}$, is a binary indicator equal to one if the algorithm tool was available to the team during their discussion of the child’s first referral $r(i)$ during the trial, and zero if the algorithm was not available. Standard errors are clustered conservatively at the household (overlapping family network) level. [^12] A household cluster is defined conservatively as any extended network of individuals who ever appeared jointly on a referral. In practice, there is no difference from clustering at the referral level (Appendix Table A.5).
@@
The effectiveness of algorithmic decision aids depends on whether workers choose to use them. Since workers were strongly encouraged but not required to consult the tool in our setting, the policy-relevant treatment effect is the intent-to-treat (ITT) effect of access to the algorithm. Using de-identified discussion notes, we ascertain that workers wrote down algorithm scores in at least 73 percent of instances when available in our sample (Table A.3). This fraction of observed and recorded scores is relatively constant across the algorithm risk score distribution (Table A.3, column 2). The recorded share is likely an underestimate: workers may have consulted the tool but forgotten to write down the score in their meeting notes. Our estimates show that workers regularly chose to consult the algorithm, even though there were no professional repercussions for not using it.
Random assignment is important for estimating unbiased causal effects of algorithm availability on outcomes. We check for evidence of random assignment with a balance test, regressing assignment status on a vector of salient pre-randomization observable characteristics. Table A.4 confirms that the algorithm was randomized successfully (p-value of the F-test is 0.74).
@@ 3.5 Outcomes: Proxies for Child Harm
We measure effects on child wellbeing using a pre-specified set of hospitalization outcomes across the entire state. These include: a child’s number of high-priority hospital admissions (emergency, urgent, or trauma-related), admissions with injury ICD (diagnosis) codes, preventable emergency room visits, visits related to substance abuse and exposure, visits with ICD codes for intentional harm (assault or self-harm), and visits with ICD codes for maltreatment, which are rare. Preventable ER visits are defined as the number of child admissions for ambulatory care sensitive conditions (ACSC, classified using Appendix 8 of Carey et al. 2017). For example, some ER visits for asthma may have been avoidable if a child had received proper preventative treatment and supervision.
@@
These categories of hospitalization have been clinically and empirically associated with child maltreatment. Maltreatment ICD codes are a measure of physician-confirmed maltreatment, but notoriously underestimate overall child maltreatment (Scott et al., 2009; Schnitzer et al., 2011). Consequently, researchers have primarily used and validated injuries as a leading proxy for maltreatment (Schnitzer et al., 2011; Vaithianathan et al., 2020). Prior research has also demonstrated a link between maltreatment and intentional injury, namely self-harm, which can often result from adverse childhood experiences such as parental neglect and sexual abuse (Liu et al. 2018, Vaithianathan et al. 2020). High-priority admissions are a related signal of potential acute harm, and avoidable ER visits are particularly associated with possible neglect and insufficient parental supervision. Substance abuse and exposure (i.e., from parents or youths’ direct abuse) are one of the primary drivers of child neglect and CPS involvement more broadly. Our outcome categories therefore cover a range of maltreatment types: physical harm, psychological harm, and neglect.
@@
Our preferred outcome is an aggregated “harm index” of pre-specified hospitalization outcomes. A high Cronbach’s alpha ($\alpha=0.85$), reflecting internal consistency across the outcomes, confirms that combining these outcomes into a composite index is meaningful. The index is the standardized average of the six pre-specified variables described above, each themselves standardized. Variables are standardized (mean 0, variance 1) using the control sample, and a lower value of the harm index is considered better for the child. This index was chosen as the most standard and agnostic way to combine these outcomes, in the absence of a pre-specified composite measure. Its advantage is to equally weigh all of the six proxies for maltreatment, regardless of each outcome’s baseline prevalence. This is in contrast to measuring the number of distinct (single-counted) hospital visits that include at least one category from the index (Row 5 in Table A.5), which in practice only picks up on the most prevalent harm index component, the number of high-priority visits (correlation $\rho=0.98$). The harm index weighs equally prevalent and less prevalent outcomes and aligns better with the idea that one of these health incidents is not per se an indication of maltreatment, but rather that having an unusually high number of visits on several of these margins at the same time is suggestive of concern. In Appendix Table A.5 we consider alternative indices: an optimally weighted index if all components are equally impacted (O’Brien, 1984), a first principal component, and two ways of counting hospital visits for harm-index-related categories. Per our pre-analysis plan, all outcomes are measured following 30 days after randomization to accommodate a period of potential mechanical increases in child receipt of medical care from CPS interactions, such as post-investigation medical care for previously untreated conditions. Results are robust to excluding the first 60 days or to including the first 30 days (Table A.5).
@@ 4 Results
Despite Child Protective Service’s primary objective of protecting children, few studies have produced causal estimates of the impact of CPS policies on child injury and other health outcomes. We link statewide hospitalization records from the trial state’s hospital association (Section 3.1) to measure the effect of algorithmic information on ground-truth outcomes. The algorithm’s score, a summary measure of family history with social services, appears to have helped workers respond more effectively to current family concerns, targeting visits to children in greater need, regardless of their past contact with CPS (see Section 5 for more detail on mechanisms). This section covers three main findings: a reduction in child harm, a reduction in child harm disparities, and a reduction in racial CPS surveillance disparities.
@@ 4.1 Impact of Algorithm Availability on Child Health and Wellbeing
Table 1 presents estimated effects of algorithm support on child wellbeing for the outcomes specified in the previous subsection, using equation (1). Column 1 of Panel A shows that tool availability reduced a child’s future harm by 0.05 standard deviations (SDs) compared to children in the control group ($p=0.046$). Effects seem attributable to both the intensive and extensive margins (see Table A.6, Column 2 for a harm index constructed using just the extensive margin). Estimates on the full sample likely underestimate our treatment effects. Children referred late in the trial are more likely to have limited or no hospital interactions prior to the trial end date, [^13] At the end of the trial, the algorithmic tool became available to all children being (re-)referred, attenuating expected treatment effects. leading to non-classical measurement error in the outcome which would bias our estimates toward zero. As a result, when we restrict to the sample of children referred in the first year of the trial, effect sizes increase to $0.06$ SD and several of the reductions in the harm components (Panel B) become significant at the 95 percent confidence level. As we did not pre-register the sample in column 2, we report estimates using the full sample throughout.
A 0.05 SD reduction in the harm index is sizable in this context. For comparison, children removed to foster care during the trial – a proxy for severe maltreatment or significant danger – had an accumulated harm index that was 0.46 standard deviations greater than those children who were not removed. The intent-to-treat effect from the algorithm is approximately 11 percent of this benchmark and – assuming our first stage does not underestimate tool use – the effect on compliers (-0.07 SD; Table A.5) is as large as 15 percent of the benchmark.
Column 3 of Panel A shows that the algorithmic tool was particularly effective at reducing extreme values of child harm (-60%, $p=0.0031$, not pre-registered): Significantly fewer children placed in the top one percent of the harm index (defined over the intervention and control groups together) when the tool was available. Accumulated harm for this top percent ranged from 2.6 to 29 standard deviations above the mean, with the median child in this group having a high priority hospital visit every three months. Relatedly, benefits of the algorithm appear to be greatest for children in the top quintile of algorithm-predicted risk scores, who may have been at greatest risk of maltreatment (Appendix Figure A.5). The majority of children with an explicit maltreatment-related hospital ICD code had a harm index in the top one percent of the harm distribution, which suggests that the algorithm may have reduced child maltreatment.
@@
Panel A: Combined Harm Index
Full
Sample
Longer
Follow-Up
Top 1% in
Harm Index
Algorithm Available -0.050** -0.062** -0.0088***
(0.025) (0.029) (0.0031)
[p = 0.046] [p = 0.032] [p = 0.004]
Randomization Controls Yes Yes Yes
Control Mean 0.000 0.000 0.015
Effect (%) -60
Design Power 0.85 0.91 0.86
Observations 4,681 3,431 4,681
Panel B: Harm Index Components
(1) (2) (3) (4) (5) (6)
@@
Algorithm Available -0.069 -0.034* -0.016 -0.034* -0.0019 -0.0052
(0.043) (0.019) (0.017) (0.018) (0.0034) (0.0036)
[p = 0.112] [p = 0.068] [p = 0.335] [p = 0.063] [p = 0.585] [p = 0.142]
Randomization Controls Yes Yes Yes Yes Yes Yes
Control Mean 0.543 0.166 0.137 0.079 0.012 0.015
Effect (%) -13 -21 -12 -44 -16 -34
Design Power 0.83 0.73 0.60 0.61 0.30 0.52
Observations 4,681 4,681 4,681 4,681 4,681 4,681
@@
Notes: This table presents estimates of the impact of algorithm availability on child injuries and other hospitalization outcomes. In Panel A, we report effects on a combined harm index, whereas in Panel B we report effects on each of the pre-registered subcomponents of the harm index before standardization. The harm index is constructed as a standardized average of the following standardized outcomes: number of high-priority admissions (emergency, urgent, and trauma), number of admissions with listed injury, number of avoidable ER visits, number of substance abuse-related visits, number of admissions with a confirmed maltreatment code, and number of visits with an intentional injury (assault, self-harm) code. A lower value of the harm index is considered better. In Panel A, beyond our main estimate in Column 1, we also report in Column 2 the effects on a sample of children with a longer follow-up (referred in the first year of the trial) so that we can observe several months of hospital data. Column 3 estimates the effect on the likelihood of being in the top one percent of the harm index. Randomization procedure controls are included throughout. We report effects as a percent of the control mean and the power to detect significant effects at the bottom of the table. Power is calculated using observed means and standard deviations in treatment and control groups, assuming 1,400 household clusters per group, an average cluster size of two, a coefficient of variation (CV) of 0.56, an intraclass correlation (ICC) of 0.4, and a significance level of 10 percent. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
Panel B of Table 1 present effects on each of the six outcomes that make up the harm index. Point estimates are negative across all six outcomes, and significant or nearly significant for the two outcomes where we have highest power ($p=0.112$ for high priority visits and $p=0.068$ for injuries). Furthermore, the point estimates are large relative to the means in the control group, ranging from -12% to -44%. In particular, the algorithmic tool reduced the average number of future injury admissions by 21 percent, or 0.03 admissions relative to a mean of 0.17. In the appendix, we report how informative the estimates of these six components would be from a Bayesian perspective (details provided in the next section), assuming as a prior that effects could have ranged between plus or minus one hundred percent (Appendix A.6). Even for outcomes where our findings are not statistically significant at conventional levels, our findings are informative for the likely range of effects. Our final pre-registered proxy for maltreatment was a previously identified list of ICD codes that are suggestive of maltreatment for a subset of child ages, following Schnitzer et al. (2011). As these are age-specific, this outcome is only available for half of our sample, and consequently cannot be included in the full sample index. We are underpowered on this margin, but the percentage effect is consistent with a halving of incidents suggestive of maltreatment (Column 1 of Appendix Table A.6).
@@
Reassuringly, we find no effects of the tool for a natural placebo group: for 11 percent of children assigned to treatment, workers tried but could not see the score due to temporary unavailability, and resulting outcomes were comparable to the control group (Table A.6, Column 7 placebo specification). We additionally present a series of robustness tests for the main effects of the algorithm on child harm and injury-specific hospital visits in Appendix Table A.5. The findings are robust to: a permutation test ($p=0.04$); to controlling for all baseline characteristics (including prior hospitalizations); excluding the first 60 days after a referral when prior injuries could be discovered during a CPS-initiated health examination; to including all days after a referral; to alternative clustering of standard errors; and to a bounding exercise for treatment non-compliance. [^14] As discussed in Appendix C, $n=80$ children in the control group had the algorithmic tool available due to changes in family structure. The findings are also robust to other ways of combining outcomes into a harm index, such as using a summation of visits from each harm index category, a summation of distinct (non-overlapping) harm index visits, using an optimally weighted index if all components are equally impacted (O’Brien, 1984), and using a first principal component. The least statistically significant effect is for the distinct count of harm visits where the point estimate suggests a marginally significant ($p=0.100$) 0.07 decrease (13% decrease). Despite its intuitive appeal, this count of harm visits turns out to be a near proxy of the number of high priority visits ($\rho=0.98$). In contrast, our preferred outcome, the harm index, combines all margins of potential harm, including those with lower prevalence, and takes into account hospitalizations with multiple causes for concern (e.g., with both injury and substance exposure codes).
The reduction in child harm does not appear to be driven by changes in willingness to seek care. Tool access did not significantly reduce the overall incidence or frequency of hospital interactions (Columns 3 and 4 of Appendix Table A.6). Our pre-registered placebo outcome of child cancer-related hospitalizations (pre-specified following Vaithianathan et al. 2020) is insignificant, though under-powered. We also find null results for better-powered, non-pregistered placebo results for a specification where workers reported not seeing a score even when assigned to treatment, and when using lower-priority hospital visits as an outcome (columns 6 and 7 of Appendix Table A.6). Consistent with our pre-analysis plan, providing algorithm support specifically decreased hospitalizations for reasons more closely associated with maltreatment. We find an imprecise reduction in charges (cannot observe negotiated bill; column 5 of Appendix Table A.6).
We do not detect significant effects of providing the algorithmic tool on the incidence or number of re-referrals and removals (Appendix Table A.7). The absence of detectable impacts on removals should not be over-interpreted, as our study is underpowered to detect even large effects (e.g., power $\approx 0.2$ to detect a 15 percent change). Moreover, the short-term impact of better-targeted investigations on removals is theoretically ambiguous. Our point estimates indicate a small possible reduction in the number of re-referrals (–5 percent), though the lack of statistical significance is more surprising for this outcome than for removals. This pattern suggests potential discrepancies between re-referral–based measures of maltreatment and hospital-based measures, reminiscent of the long-standing literature documenting differences between self-reported and substantiated CPS maltreatment (Negriff et al., 2017). Prior research has shown that the informational quality (“signal”) of CPS referrals varies by reporter type (Nadon et al., 2023). In line with this evidence, when we restrict to re-referrals from law enforcement – a reporter type with a high substantiation rate and known to be relatively reliable – we find suggestive declines consistent with our hospital-based outcomes (–12.5 percent decrease in law enforcement re-referrals, $p=0.20$). In contrast, re-referrals from less reliable reporters, such as education professionals, show no decline (+7.6 percent, $p=0.47$). Given that CPS reporting is selective and intertwined with exposure to reporters, our findings suggest that re-referral counts may be an imperfect and noisy proxy for child welfare. Outcomes external to CPS, such as hospitalization, are therefore essential for corroborating the effectiveness of CPS interventions.
Exploring Spillovers and Learning.